OpenStudy (anonymous):

Anyone can help me solve this question? [probability and statistic] The department received 4 calls every 20 minutes. Based on the time given, find the probability that the department will received not more than 3 calls.

OpenStudy (amistre64):

do you know if this is Poison based? or not ...

OpenStudy (amistre64):

and is there a table of graph of some sorts with this?

OpenStudy (anonymous):

yes i know.. i already try but answer i get dfference from lecturer given..

OpenStudy (amistre64):

fill me in on how youve tried to work it, becasue its not ringing any bells for me at the moment

OpenStudy (amistre64):

mean = 4/20 = 1/5

OpenStudy (amistre64):

\[P(x,u) = \frac{e^{-u}u^x}{x!}\] correct?

OpenStudy (amistre64):

P(0,u)+P(1,u)+P(2,u)+P(3,u)

OpenStudy (anonymous):

yes that a formula..then when it is less than 3.. we must get when P(X=0)+P(X=1)+P(X=2) right?

OpenStudy (amistre64):

yes, but this is "not more than" 3, so we can have 0,1,2, or 3

OpenStudy (anonymous):

then the answer different also.. the answer lectrer was 0.4335..

OpenStudy (amistre64):

\[\frac{e^{-.2}.2}{0!}+\frac{e^{-.2}.2}{1!}+\frac{e^{-.2}.2^2}{2!}+\frac{e^{-.2}.2^3}{3!}\] \[e^{-.2}(1+.2+.02+..00133...)\] \[e^{-.2}(1.22133...)=.9999431...\] or have a messed it up someplace?

OpenStudy (amistre64):

maybe a cumulative from 0 to 3?

OpenStudy (anonymous):

what do u mean?

OpenStudy (amistre64):

i have a poisonCDF function on my ti83, i was wondering if it would give something else but it still gives the .9999431...

OpenStudy (anonymous):

oh so that have any other formula or what?

OpenStudy (amistre64):

http://stattrek.com/probability-distributions/poisson.aspx the bottom of the page, example 1, is similar to this problem

OpenStudy (anonymous):

may i knowthe question ask not more than 3 right? can u tell me more why u take 3 also?

OpenStudy (amistre64):

because 3 is not MORE than 3

OpenStudy (anonymous):

ok i think i get it now.. thanks a lot..

OpenStudy (amistre64):

0 is not more than 3 1 is not more than 3 2 is not more than 3 3 is not more than 3 4 IS more than 3

OpenStudy (amistre64):

i think the setup we did was correct, but we have to determine if the mean we are using is accurate

OpenStudy (amistre64):

are we wanting to determine the probabiliy of the rate of x calls in 20 minutes? how do we determine the value of u?

OpenStudy (anonymous):

i think the lambda was 4,, bcozthe question ask to find not more than 3 and do not have specific time at there.. but i'm not sure..

OpenStudy (amistre64):

hmm, with 4, we do get the .43347...

OpenStudy (amistre64):

i spose it was to be interpreted as 4 per 1 time interval ...

OpenStudy (anonymous):

i think so.. but thank u for ur helping.. :)

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

thanks..u too gud luck..